### v1

# MER 2024: Semi-Supervised Learning, Noise Robustness, and Open-Vocabulary Multimodal Emotion Recognition

Abstract.
Multimodal emotion recognition is an important research topic in artificial intelligence. Over the past few decades, researchers have made remarkable progress by increasing dataset size and building more effective architectures. However, due to various reasons (such as complex environments and inaccurate labels), current systems still cannot meet the demands of practical applications. Therefore, we plan to organize a series of challenges around emotion recognition to further promote the development of this field. Last year, we launched MER2023111http://merchallenge.cn/mer2023, focusing on three topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue to organize MER2024222https://zeroqiaoba.github.io/MER2024-website. In addition to expanding the dataset size, we introduce a new track around open-vocabulary emotion recognition. The main consideration for this track is that existing datasets often fix the label space and use majority voting to enhance annotator consistency, but this process may limit the model’s ability to describe subtle emotions. In this track, we encourage participants to generate any number of labels in any category, aiming to describe the character’s emotional state as accurately as possible. Our baseline is based on MERTools [^3] https://github.com/zeroQiaoba/MERTools and the code is available at: https://github.com/zeroQiaoba/MERTools/tree/master/MER2024.
MER 2024, multimodal emotion recognition, semi-supervised learning, noise robustness, open-vocabulary emotion recognition
copyright: rightsretainedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32st ACM International Conference on Multimedia; October 28-November 1, 2024; Melbourne, Australiabooktitle: Proceedings of the 32st ACM International Conference on Multimedia (MM ’24), October 28-November 1, 2024, Melbourne, Australiaprice: 15.00ccs: Computing methodologies Neural networksccs: Computing methodologies Artifcial intelligenceccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing
1. Introduction
Multimodal emotion recognition plays an important role in human-computer interaction. With the development of deep learning, researchers have made significant progress in this task. However, it still cannot be well addressed, and the emotion recognition performance still cannot meet the requirements of practical applications (lian2024merbench, ). To promote the development of this field, last year, we launched MER2023 (lian2023mer, ), focusing on three interesting topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue the latter two tracks and add a new track around open-vocabulary emotion recognition, aiming to describe the character’s emotional state as accurately as possible.
First, it is difficult to collect large amounts of samples with emotion labels. On the one hand, the collected samples usually have no emotion (i.e. neutral) (poria2019meld, ). On the other hand, emotions have inherent subjectivity. To increase annotation consistency, researchers often hire multiple annotators and adopt majority voting (li2017reliable, ), significantly increasing labor costs. To address the problem of data sparseness, researchers revolve around unsupervised or semi-supervised learning and use unlabeled data during training. Furthermore, MERBench (lian2024merbench, ) points out the necessity of using unlabeled data from the same domain as labeled data. Therefore, we provide a large number of human-centric unlabeled videos in MER2024 and encourage participants to explore more effective unsupervised or semi-supervised learning strategies for better performance.
Meanwhile, noise generally exists in the video. It is hard to guarantee that every video is free of any audio noise and each frame is in high-resolution. To improve noise robustness, researchers have carried out a series of works on emotion recognition under noise conditions (zhao2021missing, ). However, there lacks a benchmark dataset to fairly compare different strategies. Therefore, we organize a track around noise robustness. Although there are many types of noise, we consider the two most common ones: audio additive noise and image blur noise. We encourage participants to exploit data augmentation (hazarika2022analyzing, ) or other effective techniques (sun2023efficient, ; lian2023gcnet, ) to improve the noise robustness of emotion recognition systems.
Furthermore, emotions are subjective and ambiguous. To increase annotation consistency, existing datasets typically limit the label space to a few discrete labels, employ multiple annotators, and use majority voting to select the most likely label. However, this process may cause some correct but non-candidate or non-majority labels to be ignored, resulting in inaccurate annotations. To this end, we introduce a new track around open-vocabulary emotion recognition. We encourage participants to generate any number of labels in any category, trying to describe emotions accurately.
To sum up, MER2024 contains three tracks: MER-SEMI, MER-NOISE, and MER-OV. In MER-SEMI, we encourage participants to use unlabeled data for better performance; in MER-NOISE, we focus on the systems’ noise robustness; in MER-OV, we ask participants to describe the character’s emotional state as accurately as possible. This challenge provides participants with a platform to fairly compare the performance of different methods. In the remainder of this paper, we introduce the dataset, baseline systems, evaluation metrics, and experimental results for each track.
Figure 1. Dataset construction pipeline for MER2024.
2. Challenge Dataset
MER2024 is an extended version of MER2023 (lian2023mer, ). Figure 1 shows the dataset construction pipeline. Specifically, MER2023 contains four subsets: Train&\&&Val, MER-MULTI, MER-NOISE, and MER-SEMI. In the last subset, besides labeled samples, it also contains a large amount of unlabeled data. In MER2024, we merge all labeled samples and get the updated Train&\&&Val. Meanwhile, we collect more unlabeled data and select some unlabeled samples for annotation, getting MER-SEMI and MER-NOISE. For MER-OV, we select 332 samples from Train&\&&Val and rely on LLMs to extract open-vocabulary labels from emotion-related descriptions (see Section 2.3 for more details). Table 1 provides statistics for MER2023 and MER2024.
2.1. Data Collection
As shown in Figure 1, MER2024 increases the number of unlabeled samples. The data collection process is borrowed from MER2023 and includes two key steps: video cutting and video filtering.
Video Cutting. To simulate real scenarios, we collect raw video from movies and TV series. However, the raw video is often long and have multiple characters. It is impossible to assign an emotion label to the entire video. Therefore, we need to split the raw video into video clips. In this process, we require that the video clips should contain relatively complete content from the same speaker, which is beneficial to understanding the emotional state of the character. Considering that the timestamps in subtitles are generally accurate, for videos with subtitles, we directly use these timestamps for video segmentation. For videos without subtitles, we use the voice activity detection toolkit, Silero VAD [^4] https://github.com/snakers4/silero-vad, to segment the video and the speaker identification toolkit, Deep Speaker [^5] https://github.com/philipperemy/deep-speaker, to merge the consecutive clips if they are likely to be from the same speaker.
Video Filtering. In the previous step, we aim to generate video clips containing relatively complete content from the same speaker. However, it only guarantees that the audio is from the same character, not that there is one character in the visual frames. Therefore, we further filter these video clips. Specifically, we use the face detection toolkit, YuNet [^6] https://github.com/ShiqiYu/libfacedetection, to ensure that most frames contain only one face. Then, we use the face recognition toolkit, face.evoLVe [^7] https://github.com/ZhaoJ9014/face.evoLVe, to ensure that these faces belong to the same person. Additionally, the length of video clips is also important. Video clips that are too short generally cannot express the complete meaning of the character; video clips that are too long may contain emotional changes, making it difficult to accurately describe the character’s emotional state with one label. Therefore, we perform an additional filtering operation, only selecting video clips between 2$\sim$16 seconds.
Table 1. Statistics for MER2023 and MER2024. Here, “A/B” means that there are B samples in the subset and we only annotate A samples for performance evaluation.
MER2023
# of samples
(labeled/whole)
MER2024
# of samples
(labeled/whole)
Train&\&&Val 3373 Train&\&&Val 5030
MER-MULTI 411 MER-SEMI 1169/115595
MER-NOISE 412 MER-NOISE 1170/115595
MER-SEMI 834/73982 MER-OV 322
2.2. MER-SEMI and MER-NOISE
Annotating all unlabeled data requires significant labor costs. To reduce the cost, we only select samples with a high probability of explicit emotions for labeling. Specifically, we use the top-16 prediction results in last year’s challenge and calculate the proportion of primary labels as the basis for selection. For example, if a sample is predicted as happy 10 times and neutral 6 times, its score is $v=10/16$. Then, taking into account the score and class balance, we select a total of 6,000 samples for labeling. During the annotation process, we hire 5 annotators and only select samples in which at least 4 annotators assign the same label, resulting in 2,339 labeled samples. All these samples are equally divided into two parts, one for MER-SEMI and the other for MER-NOISE.
In MER-NOISE, we need to add noise to the video. This paper considers two types of noise: audio additive noise and image blur noise. For the audio modality, we select noise from the MUSAN dataset (snyder2015musan, ), which contains three subsets: music, noise, and speech. The noise in the first two subsets may convey emotions, thus affecting the emotion of the raw audio. For example, rain and thunder may lead to a negative sentiment, while a pleasant breeze may generate a positive sentiment. Therefore, we only use the noise in the last subset. In this process, we randomly select a speech-to-noise ratio (SNR) between 5dBsimilar-$\sim$10dB and randomly select noise from the speech subset. For the visual modality, image blur is a common noise. To generate blurry images, we downsample the image to lose some details, and then upsample the low-resolution image to keep the size unchanged. In this process, we randomly select a downsampling factor from $r=\{1,2,4\}$.
2.3. MER-OV
MER-OV focuses on open-vocabulary emotion recognition. Unlike MER-SEMI and MER-NOISE which only predict the most likely emotion within a fixed label space, MER-OV needs to predict any number of emotions in any category, trying to describe the character’s emotional state as accurately as possible. Previously, we made an initial attempt on this task (lian2023explainable, ). Specifically, we first annotate emotion-related acoustic and visual clues. Then, we rely on the LLMs’ reasoning ability and use these clues to disambiguate video subtitles. This process can generate descriptions containing more subtle emotional labels. Finally, we use ChatGPT to extract all labels using the following prompt:
Please assume the role of an expert in the field of emotions. We provide clues that may be related to the emotions of the characters. Based on the provided clues, please identify the emotional states of the main characters. Please separate different emotional categories with commas and output only the clearly identifiable emotional categories in list format. If none are identified, please output an empty list.
Due to the high annotation cost, we only select 332 samples for annotation (lian2023explainable, ). Through this process, we can generate an average of 2.92 labels per sample. We encourage participants to use MLLMs directly or to further conduct instruction fine-tuning with our annotated samples to solve this task.
Table 2. Prompt for generating emotion-related descriptions using MLLMs.
Models Prompt
Audio LLM Subtitle content of the audio: {subtitle}; As an expert in the field of emotions, please focus on the acoustic information and subtitle content in the audio to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the audio.
Video LLM Subtitle content of the video: {subtitle}; As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
Audio-Video LLM Subtitle content of the video: {subtitle} As an expert in the field of emotions, please focus on the facial expressions, body movements, environment, acoustic information, subtitle content, etc., in the video to discern clues related to the emotions of the individual. Please provide a detailed description and ultimately predict the emotional state of the individual in the video.
2.4. Challenge Protocol
To download the dataset, participants should fill out an End User License Agreement (EULA) [^8] https://drive.google.com/file/d/1I0ZuPzL96W4Ow3KdF3N6D-mDhS5qJc-x. It requires participants to use this dataset only for academic research and not to edit or upload it to the Internet. For MER-SEMI and MER-NOISE, each team should submit the most likely discrete label among the 6 candidate labels (i.e., worried, happy, neutral, angry, surprised, and sad). For MER-OV, each team can submit any number of labels in any category. For all tracks, participants should predict results for all 115,595 unlabeled data, although we only evaluate a small subset. This requires participants to focus on the generalization ability and develop systems that do not require a large amount of inference time. Additionally, it is worth noting that participants cannot manually annotate the samples in MER2024. We will ask them to submit their code for further checking. Finally, each team should submit a paper describing their method, otherwise, the submitted results will not be included in the final leaderboard.
3. Baselines and Evaluation Metrics
We first introduce the baselines of these tracks. For MER-SEMI and MER-NOISE, we establish the baseline based on MERTools [^9] https://github.com/zeroQiaoba/MERTools. For MER-OV, we use MLLMs as the baselines since these models can solve various multimodal tasks. Then, we illustrate implementation details and evaluation metrics for each track.
3.1. MER-SEMI and MER-NOISE
Feature selection and fusion are crucial to building an effective emotion recognition system. Regarding feature selection, we adopt the recommendations of MERBench (lian2024merbench, ). Language compatibility is important for both lexical and acoustic features. Therefore, we mainly choose encoders trained in Chinese corpora. For visual features, domain compatibility is also important. Therefore, in addition to encoders trained on action or caption datasets, we also select encoders trained on human-centric videos. Table 6 lists model cards for some representative unimodal encoders.
Regarding fusion strategies, MERBench points out that the attention mechanism can achieve relatively good performance among all algorithms (lian2024merbench, ). The reason lies in that the labeled samples are limited in emotion recognition. A complex architecture can easily lead to the overfitting problem, affecting the model’s generalization ability on unseen data. Specifically, assume that acoustic, lexical, and visual features are $f_{a}\in\mathbb{R}^{d_{a}}$, $f_{l}\in\mathbb{R}^{d_{l}}$, $f_{v}\in\mathbb{R}^{d_{v}}$, respectively. During the fusion process, we first map these features to a fixed dimension $d$:
(1) $h_{m}=\mbox{ReLU}\left(f_{m}W_{m}+b_{m}\right),m\in\{a,l,v\},$
where $W_{m}\in\mathbb{R}^{d_{m}\times d}$ and $b_{m}\in\mathbb{R}^{d}$ are trainable parameters. Then, we calculate the attention score for each modality:
(2) $h=\mbox{Concat}\left(h_{a},h_{l},h_{v}\right),$
(3) $\alpha=\mbox{softmax}\left(h^{T}W_{\alpha}+b_{\alpha}\right),$
where $W_{\alpha}\in\mathbb{R}^{d\times 1}$ and $b_{\alpha}\in\mathbb{R}^{1}$ are trainable parameters. Here, $h\in\mathbb{R}^{d\times 3}$ and $\alpha\in\mathbb{R}^{3\times 1}$. The fused multimodal feature $z=h\alpha\in\mathbb{R}^{d}$ is used for subsequent emotion recognition.
3.2. MER-OV
Open-vocabulary emotion recognition is a new task that lacks large-scale training data. Considering that MLLMs can handle various multimodal tasks, we choose MLLMs as the baselines. Meanwhile, temporal information is important for understanding the character’s emotions. Therefore, we select MLLMs that support at least audio or video input, including video LLMs (such as VideoChat (li2023videochat, )), audio LLMs (such as SALMONN (tang2023salmonn, )), and audio-video LLMs (such as PandaGPT (su2023pandagpt, )). In practice, we first use the prompt in Table 2 to generate emotion-relevant descriptions that contain multi-faceted and multi-modal clues. During inference, we add subtitles as additional input. Then, we use the prompt in Section 2.3 to extract all emotion labels in an open-vocabulary manner.
3.3. Implementation Details
For MER-OV, we directly use pretrained MLLMs. We rely on GPT-3.5 (“gpt-3.5-turbo-16k-0613”) to generate emotions in an open-vocabulary manner. For MER-SEMI and MER-NOISE, we use the attention mechanism for multimodal fusion. This process involves one hyper-parameter, i.e., the dimension of hidden representation $d$, and we choose $d$ from $\{64,128,256\}$. During training, we use the Adam optimizer and choose the learning rate from $\{10^{-3},10^{-4}\}$. To alleviate the overfitting problem, we also use Dropout and select the rate from $\{0.2,0.3,0.4,0.5\}$. Therefore, a total of 3 hyper-parameters need to be adjusted. To find the optimal hyper-parameters, we randomly select 50 parameter combinations and choose the model that performed best in five-fold cross-validation. Then, we report its average result and standard deviation.
Table 3. Unimodal results. In addition to five-fold cross-validation results on Train&\&&Val, we also report results on MER-SEMI and MER-NOISE. These features are ordered by the average WAF score on Train&\&&Val, MER-SEMI, and MER-NOISE. The values in gray columns are used as the final ranking metric for these tracks.
Feature Train&\&&Val MER-SEMI MER-NOISE Average
WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ )
Visual Modality
VideoMAE-base (tong2022videomae, ) 52.86$\pm$0.23 53.27$\pm$0.21 47.37$\pm$0.41 52.04$\pm$0.16 48.53$\pm$0.39 51.81$\pm$0.26 49.59 52.37
EmoNet (kahou2016emonets, ) 51.76$\pm$0.31 53.18$\pm$0.10 52.45$\pm$0.18 54.94$\pm$0.15 51.22$\pm$0.21 53.56$\pm$0.43 51.81 53.90
VideoMAE-large (tong2022videomae, ) 57.04$\pm$0.25 57.49$\pm$0.30 51.05$\pm$0.39 55.73$\pm$0.41 50.81$\pm$0.28 55.22$\pm$0.62 52.97 56.15
DINOv2-large (oquab2023dinov2, ) 58.44$\pm$0.12 59.57$\pm$0.15 53.41$\pm$0.26 57.47$\pm$0.22 52.37$\pm$0.31 56.10$\pm$0.14 54.74 57.71
SENet-FER2013 (hu2018squeeze, ) 57.67$\pm$0.10 58.79$\pm$0.16 54.62$\pm$0.16 56.36$\pm$0.24 52.31$\pm$0.24 54.79$\pm$0.16 54.87 56.65
ResNet-FER2013 (he2016deep, ) 58.73$\pm$0.30 59.66$\pm$0.16 53.07$\pm$0.20 55.13$\pm$0.27 53.26$\pm$0.29 55.54$\pm$0.20 55.02 56.78
MANet-RAFDB (zhao2021learning, ) 59.91$\pm$0.12 61.10$\pm$0.11 56.59$\pm$0.31 58.87$\pm$0.43 55.15$\pm$0.30 57.49$\pm$0.23 57.22 59.15
EVA-02-base (fang2023eva, ) 61.41$\pm$0.20 62.28$\pm$0.22 55.25$\pm$0.21 58.80$\pm$0.24 55.36$\pm$0.14 58.14$\pm$0.15 57.34 59.74
CLIP-base (radford2021learning, ) 61.74$\pm$0.18 62.56$\pm$0.20 57.16$\pm$0.14 60.69$\pm$0.24 57.43$\pm$0.16 61.25$\pm$0.19 58.78 61.50
VideoMAE-base (VoxCeleb2) (chung2018voxceleb2, ) 63.33$\pm$0.13 63.84$\pm$0.14 57.98$\pm$0.20 61.56$\pm$0.23 57.18$\pm$0.28 60.06$\pm$0.19 59.50 61.82
VideoMAE-base (MER2023) (lian2023mer, ) 64.50$\pm$0.17 64.93$\pm$0.19 58.34$\pm$0.32 62.10$\pm$0.35 57.32$\pm$0.20 61.07$\pm$0.31 60.05 62.70
CLIP-large (radford2021learning, ) 66.66$\pm$0.12 67.17$\pm$0.12 63.27$\pm$0.29 65.45$\pm$0.30 60.04$\pm$0.19 62.78$\pm$0.13 63.32 65.13
Acoustic Modality
eGeMAPS (eyben2016geneva, ) 39.68$\pm$0.53 42.88$\pm$0.38 30.27$\pm$0.58 33.85$\pm$0.79 28.92$\pm$0.39 32.05$\pm$0.43 32.96 36.26
VGGish (hershey2017cnn, ) 48.60$\pm$0.10 50.20$\pm$0.09 45.44$\pm$0.37 48.52$\pm$0.23 40.70$\pm$0.31 43.69$\pm$0.23 44.91 47.47
Whisper-base (radford2023robust, ) 56.65$\pm$0.13 57.08$\pm$0.24 59.70$\pm$0.16 60.60$\pm$0.12 41.26$\pm$0.41 42.52$\pm$0.23 52.54 53.40
emotion2vec (ma2023emotion2vec, ) 56.08$\pm$0.09 56.48$\pm$0.03 58.98$\pm$0.33 59.68$\pm$0.26 45.66$\pm$0.31 46.97$\pm$0.31 53.57 54.38
Whisper-large (radford2023robust, ) 63.23$\pm$0.20 63.27$\pm$0.20 69.24$\pm$0.87 70.62$\pm$0.62 53.26$\pm$0.86 55.34$\pm$0.57 61.91 63.08
wav2vec 2.0-base (baevski2020wav2vec, ) 64.89$\pm$0.26 65.14$\pm$0.30 68.12$\pm$0.34 69.29$\pm$0.25 53.91$\pm$0.14 56.87$\pm$0.29 62.31 63.76
wav2vec 2.0-large (baevski2020wav2vec, ) 65.50$\pm$0.29 65.83$\pm$0.27 68.12$\pm$0.30 70.06$\pm$0.22 56.41$\pm$0.73 59.12$\pm$0.63 63.35 65.00
HUBERT-base (hsu2021hubert, ) 69.26$\pm$0.13 69.43$\pm$0.15 78.70$\pm$0.20 79.28$\pm$0.20 59.55$\pm$0.70 59.76$\pm$0.82 69.17 69.49
HUBERT-large (hsu2021hubert, ) 73.02$\pm$0.13 73.10$\pm$0.15 83.42$\pm$0.37 84.00$\pm$0.32 73.21$\pm$0.36 74.03$\pm$0.31 76.55 77.05
Lexical Modality
XLNet-base (yang2019xlnet, ) 48.59$\pm$0.30 48.95$\pm$0.29 49.13$\pm$0.07 48.64$\pm$0.04 46.14$\pm$0.16 46.44$\pm$0.10 47.96 48.01
ELECTRA-large (clark2020electra, ) 50.45$\pm$0.15 50.83$\pm$0.09 52.07$\pm$0.16 51.81$\pm$0.12 47.34$\pm$0.18 47.14$\pm$0.12 49.95 49.93
MOSS-7B 51.25$\pm$0.23 51.64$\pm$0.19 51.62$\pm$0.22 51.64$\pm$0.14 47.87$\pm$0.19 48.42$\pm$0.29 50.25 50.57
PERT-large (cui2022pert, ) 50.36$\pm$0.23 50.69$\pm$0.10 53.47$\pm$0.17 53.26$\pm$0.13 48.76$\pm$0.23 48.83$\pm$0.22 50.86 50.93
PERT-base (cui2022pert, ) 50.26$\pm$0.08 50.62$\pm$0.12 53.64$\pm$0.21 53.50$\pm$0.18 49.29$\pm$0.18 49.44$\pm$0.18 51.06 51.19
LERT-large (cui2022lert, ) 52.22$\pm$0.18 52.38$\pm$0.19 52.40$\pm$0.20 52.18$\pm$0.28 49.17$\pm$0.44 49.14$\pm$0.49 51.26 51.24
ELECTRA-base (clark2020electra, ) 50.98$\pm$0.07 51.30$\pm$0.07 54.50$\pm$0.33 54.35$\pm$0.30 49.35$\pm$0.15 49.32$\pm$0.16 51.61 51.65
LERT-base (cui2022lert, ) 52.37$\pm$0.13 52.72$\pm$0.10 54.84$\pm$0.21 54.62$\pm$0.20 47.98$\pm$0.07 48.15$\pm$0.15 51.73 51.83
RoBERTa-base (liu2019roberta, ) 51.84$\pm$0.22 52.24$\pm$0.18 54.37$\pm$0.24 53.99$\pm$0.34 49.49$\pm$0.51 49.56$\pm$0.41 51.90 51.93
MacBERT-base (cui2020revisiting, ) 51.40$\pm$0.13 51.75$\pm$0.13 55.11$\pm$0.29 54.74$\pm$0.37 49.68$\pm$0.24 49.74$\pm$0.25 52.06 52.08
RoBERTa-large (liu2019roberta, ) 52.66$\pm$0.18 52.92$\pm$0.13 55.14$\pm$0.31 54.88$\pm$0.31 49.06$\pm$0.21 49.23$\pm$0.28 52.29 52.34
ChatGLM2-6B (du2022glm, ) 53.04$\pm$0.23 53.28$\pm$0.22 55.52$\pm$0.56 55.06$\pm$0.63 50.39$\pm$0.43 50.40$\pm$0.40 52.98 52.91
MacBERT-large (cui2020revisiting, ) 52.19$\pm$0.14 52.47$\pm$0.12 56.99$\pm$0.20 56.81$\pm$0.31 50.24$\pm$0.20 50.34$\pm$0.23 53.14 53.21
BLOOM-7B (workshop2022bloom, ) 53.13$\pm$0.24 53.30$\pm$0.24 56.01$\pm$0.36 55.99$\pm$0.40 50.38$\pm$0.21 50.58$\pm$0.13 53.18 53.29
Baichuan-13B (yang2023baichuan, ) 54.86$\pm$0.12 55.11$\pm$0.06 56.63$\pm$0.18 56.17$\pm$0.34 52.01$\pm$0.30 51.90$\pm$0.21 54.50 54.39
Table 4. Multimodal results. We report both bimodal and trimodal results. Here, “A”, “V”, “T” represent acoustic, visual, and textual modalities, respectively. “TopN” means that we select the top-N features for each modality. Their ranking is based on the average WAF in Table 3.
# Top Train&\&&Val MER-SEMI MER-NOISE Average
WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ ) WAF ()($\uparrow$)( ↑ ) ACC ()($\uparrow$)( ↑ )
A+V
Top1 78.86$\pm$0.17 78.98$\pm$0.13 84.07$\pm$0.17 84.60$\pm$0.16 78.14$\pm$0.31 79.43$\pm$0.22 80.36 81.00
Top2 78.92$\pm$0.04 78.97$\pm$0.11 84.20$\pm$0.15 84.73$\pm$0.13 77.33$\pm$0.14 78.55$\pm$0.18 80.15 80.75
Top3 79.16$\pm$0.11 79.18$\pm$0.16 84.09$\pm$0.15 84.67$\pm$0.15 77.26$\pm$0.38 78.70$\pm$0.33 80.17 80.85
Top4 79.11$\pm$0.08 79.23$\pm$0.09 84.12$\pm$0.11 84.72$\pm$0.11 77.16$\pm$0.15 78.55$\pm$0.15 80.13 80.83
Top5 79.17$\pm$0.07 79.25$\pm$0.10 84.38$\pm$0.23 84.92$\pm$0.22 77.05$\pm$0.17 78.46$\pm$0.32 80.20 80.88
Top6 79.12$\pm$0.09 79.25$\pm$0.12 84.14$\pm$0.20 84.72$\pm$0.19 77.54$\pm$0.34 78.78$\pm$0.22 80.27 80.91
A+T
Top1 73.67$\pm$0.13 73.82$\pm$0.09 84.06$\pm$0.35 84.57$\pm$0.32 74.74$\pm$0.26 75.61$\pm$0.23 77.49 78.00
Top2 73.75$\pm$0.19 73.84$\pm$0.17 84.50$\pm$0.36 85.01$\pm$0.35 73.38$\pm$0.46 74.25$\pm$0.59 77.21 77.70
Top3 73.97$\pm$0.11 73.97$\pm$0.05 84.64$\pm$0.16 85.10$\pm$0.17 73.67$\pm$0.39 74.32$\pm$0.39 77.42 77.80
Top4 73.82$\pm$0.22 73.87$\pm$0.22 84.37$\pm$0.57 84.85$\pm$0.54 73.90$\pm$0.30 74.36$\pm$0.29 77.36 77.69
Top5 73.73$\pm$0.19 73.80$\pm$0.18 84.08$\pm$0.30 84.64$\pm$0.27 73.53$\pm$0.14 74.20$\pm$0.15 77.11 77.55
Top6 73.68$\pm$0.12 73.70$\pm$0.11 84.31$\pm$0.36 84.80$\pm$0.31 73.99$\pm$0.49 74.62$\pm$0.40 77.32 77.71
V+T
Top1 72.28$\pm$0.16 72.47$\pm$0.11 77.34$\pm$0.30 77.74$\pm$0.28 71.98$\pm$0.42 73.38$\pm$0.31 73.87 74.53
Top2 73.88$\pm$0.08 74.03$\pm$0.07 77.17$\pm$0.43 77.89$\pm$0.39 71.99$\pm$0.21 73.72$\pm$0.11 74.34 75.21
Top3 74.29$\pm$0.09 74.35$\pm$0.10 77.98$\pm$0.37 78.60$\pm$0.39 72.61$\pm$0.34 74.26$\pm$0.35 74.96 75.74
Top4 74.06$\pm$0.12 74.22$\pm$0.09 78.31$\pm$0.34 79.08$\pm$0.33 72.51$\pm$0.18 74.26$\pm$0.13 74.96 75.85
Top5 74.04$\pm$0.10 74.20$\pm$0.10 77.78$\pm$0.23 78.50$\pm$0.25 73.12$\pm$0.41 74.81$\pm$0.33 74.98 75.84
Top6 74.33$\pm$0.17 74.48$\pm$0.21 77.83$\pm$0.26 78.53$\pm$0.26 73.03$\pm$0.36 74.70$\pm$0.27 75.07 75.90
A+V+T
Top1 79.31$\pm$0.03 79.40$\pm$0.04 86.73$\pm$0.13 87.09$\pm$0.16 79.47$\pm$0.29 80.67$\pm$0.19 81.84 82.39
Top2 80.05$\pm$0.06 80.07$\pm$0.09 85.94$\pm$0.25 86.44$\pm$0.24 79.11$\pm$0.33 80.33$\pm$0.29 81.70 82.28
Top3 80.28$\pm$0.16 80.33$\pm$0.12 86.27$\pm$0.22 86.74$\pm$0.18 78.75$\pm$0.20 80.17$\pm$0.19 81.77 82.41
Top4 80.02$\pm$0.11 80.10$\pm$0.11 86.19$\pm$0.14 86.67$\pm$0.16 79.62$\pm$0.27 80.82$\pm$0.28 81.94 82.53
Top5 80.17$\pm$0.13 80.21$\pm$0.12 85.93$\pm$0.15 86.45$\pm$0.17 78.91$\pm$0.28 80.36$\pm$0.22 81.67 82.34
Top6 80.34$\pm$0.10 80.43$\pm$0.12 86.32$\pm$0.22 86.86$\pm$0.19 79.24$\pm$0.40 80.33$\pm$0.28 81.97 82.54
3.4. Evaluation Metrics
For MER-SEMI and MER-NOISE, we choose two widely used metrics in emotion recognition: accuracy and weighted average F-score (WAF) (lian2021ctnet, ). Considering the inherent class imbalance, we choose WAF as the final ranking. For MER-OV, we draw on our previous work (lian2023explainable, ), in which we extend traditional classification metrics (i.e., accuracy and recall) and define set-level metrics. Specifically, assume that the true label set is $\{y_{i}\}_{i=1}^{M}$ and the predicted label set is $\{\hat{y}_{i}\}_{i=1}^{N}$, where $M$ and $N$ are the number of labels. Considering that there may be labels with different expressions but the same meaning, we first use ChatGPT to group all labels:
Please assume the role of an expert in the field of emotions. We provide a set of emotions. Please group the emotions, with each group containing synonyms or consistent emotional terms. Directly output the results, with each group in list format.
Following that, we map these labels to their group IDs, where the mapping function is defined as $G(\cdot)$:
(4) ${Y}=\{G(x)|x\in\{{y}_{i}\}_{i=1}^{M}\},$
(5) $\hat{Y}=\{G(x)|x\in\{\hat{y}_{i}\}_{i=1}^{N}\}.$
The set-level accuracy and recall are defined as follows:
(6) $\mbox{Accuracy}_{{s}}=\frac{|{Y}\cap\hat{{Y}}|}{|\hat{{Y}}|},$
(7) $\mbox{Recall}_{{s}}=\frac{|{Y}\cap\hat{{Y}}|}{|{Y}|}.$
For MER-OV, we choose the average between $\mbox{Accuracy}_{{s}}$ and $\mbox{Recall}_{{s}}$ as the final ranking:
(8) $\mbox{Avg}=\frac{\mbox{Accuracy}_{{s}}+\mbox{Recall}_{{s}}}{2}.$
To obtain more accurate open-vocabulary labels, we first extract labels using the approach in Section 2.3. Then, we merge these labels with the predictions submitted by participants. Finally, we manually check all labels to obtain the final ground truth.
4. Results and Discussion
In this section, we report baseline results for three tracks. For MER-SEMI and MER-NOISE, we report both unimodal and multimodal results. For MER-OV, we report the results of various MLLMs.
4.1. MER-SEMI and MER-NOISE
Table 3 presents the unimodal results. From this table, we observe that the model that performs well on Train&\&&Val generally performs well on MER-SEMI and MER-NOISE. These results indicate that the model trained on our dataset has a good generalization ability. Among all modalities, the acoustic encoder can achieve the best unimodal result, suggesting that the samples in MER2024 mainly rely on the acoustic modality to express emotions.
For the visual modality, unsupervised or semi-supervised models generally perform better than supervised models (e.g., CLIP-large performs better than SENet-FER2013). This phenomenon shows that unsupervised or semi-supervised models can capture universal representations, which are beneficial for downstream tasks. Meanwhile, previous works point out that domain compatibility is important for the visual encoder (lian2024merbench, ). Therefore, we further train VideoMAE-based on in-domain corpora (such as VoxCeleb2 and MER2023). We observe that this process leads to significant performance improvements compared to the raw VideoMAE-based. Therefore, we recommend that participants try to train other visual encoders on in-domain corpora.
For acoustic and lexical modalities, based on the recommendations of MERBench (lian2024merbench, ), we mainly select the models trained on the Chinese corpus. From Table 3, we observe that unsupervised or semi-supervised acoustic encoders (e.g., HUBERT-large) generally perform better than traditional acoustic encoders (e.g., eGeMAPS), which is consistent with findings in the visual modality. Furthermore, LLMs often perform better than pretrained language models (PLMs). These results show that by increasing training data and model size, we can build more powerful lexical encoders and achieve better performance on emotion recognition.
Table 4 reports multimodal results. In this table, we choose “TopN” features and use the attention mechanism for multimodal fusion. “TopN” means that we select the top-$N$ features for each modality, and their ranking is based on the average WAF in Table 3. Experimental results show that different modality combinations prefer distinct $N$. Therefore, we should adjust this parameter carefully. Meanwhile, we can achieve the best result when considering trimodal information, reflecting the effectiveness of multimodal fusion. In terms of WAF scores, our baseline system can reach 86.73 in MER-SEMI and 79.62 in MER-NOISE.
Table 5. Performance of different methods on MER-OV (reported in our previous work (lian2023explainable, )). The values in the gray column are used for the final ranking. Here, “L”, “V”, and “A” indicate whether lexical, visual, and acoustic information is used in the emotion recognition process.
Model L V A Avg $\mbox{Accuracy}_{\mbox{s}}$ $\mbox{Recall}_{\mbox{s}}$
Empty $\times$ $\times$ $\times$ 0.0 0.0 0.0
Random $\times$ $\times$ $\times$ 19.0 24.1 13.8
Valley (luo2023valley, ) $\surd$ $\surd$ $\times$ 11.1 12.7 9.6
Otter (li2023otter, ) $\surd$ $\surd$ $\times$ 24.5 27.6 21.4
PandaGPT (su2023pandagpt, ) $\surd$ $\surd$ $\surd$ 25.3 26.5 24.2
Video-LLaMA (zhang2023video, ) $\surd$ $\surd$ $\times$ 28.4 27.3 29.4
VideoChat2 (li2024mvbench, ) $\surd$ $\surd$ $\times$ 32.9 36.8 28.9
Video-ChatGPT (maaz2023video, ) $\surd$ $\surd$ $\times$ 35.3 39.0 31.7
VideoChat (li2023videochat, ) $\surd$ $\surd$ $\times$ 35.9 34.3 37.5
SALMONN (tang2023salmonn, ) $\surd$ $\times$ $\surd$ 36.0 34.3 37.8
mPLUG-Owl (ye2023mplugowl, ) $\surd$ $\surd$ $\times$ 39.8 38.8 40.7
Qwen-Audio (chu2023qwen, ) $\surd$ $\times$ $\surd$ 46.0 54.0 38.0
GPT-4V (openai2023gpt4v, ) $\surd$ $\surd$ $\times$ 56.0 53.2 58.9
4.2. MER-OV
In this section, we report the performance of different methods on this track. Besides MLLMs, we introduce two heuristic baselines: Empty and Random. For the former, we do not assign any label. For the latter, we randomly select a label from six candidate categories (i.e., worried, happy, neutral, angry, surprised, and sad). Considering that the emotion labels extracted in Section 2.3 are relatively accurate, we use them as the ground truth. Table 5 shows the performance of different methods. Experimental results demonstrate that MLLMs generally outperform heuristic baselines, showcasing that MLLMs can handle this task to some extent. However, these models still cannot address this task perfectly. Meanwhile, audio LLMs generally perform better than video LLMs. The reason is that the samples in our dataset mainly convey emotions through the acoustic modality, which is consistent with the findings in Table 3.
5. Conclusions
This paper introduces a multimodal emotion recognition challenge, MER2024, which consists of three tracks: MER-SEMI, MER-NOISE, and MER-OV. This challenge is an extended version of MER2023 with more labeled and unlabeled samples. Meanwhile, it introduces a new task called open-vocabulary emotion recognition, which aims to generate any number of labels in any category to describe the character’s emotional state as accurately as possible. In this paper, we introduce the datasets, establish the baselines, define the evaluation metrics, and report the baseline results for three tracks. Meanwhile, we open-source the code to ensure reproducibility. We hope that the MER series of challenges can provide researchers with a platform to evaluate their emotion recognition systems and prompt the application of this technology in real-life scenarios.
6. Acknowledgements
We would like to thank the members of data chairs Bin Liu (Institute of Automation, CAS), Rui Liu (Inner Mongolia University), and Kele Xu (National University of Defense Technology). Meanwhile, we appreciate the program committee for their valuable support: Xiaobai Li (Zhejiang University), Zixing Zhang (Hunan University), Jianfei Yu (Nanjing University of Science and Technology), Ya Li (Beijing University of Posts and Telecommunications), Mengyue Wu (Shanghai Jiao Tong University), Jing Han (University of Cambridge), Jinming Zhao (Renmin University of China), Mingyue Niu (Yanshan University), Yongwei Li (Institute of Automation, CAS), and Licai Sun (University of Chinese Academy of Sciences). This work is supported by the National Natural Science Foundation of China (No. 62201572, No. 62276259, No. U21B2010, No. 62271083, No. 62306316, No. 62322120).
Appendix A Model Cards
MERBench [1] provides some guidelines for feature selection. For the lexical and acoustic modalities, language matching is important. Therefore, we mainly choose encoders trained in Chinese corpora. For the visual modality, domain matching is important. Therefore, we provide two visual encoders trained on human-centric videos (i.e., MER2023-SEMI and VoxCeleb2). Table 6 provides model cards for some feature extractors.
Table 6. Model cards for some feature extractors. Here, “CH”, “EN”, and “MULTI” are the abbreviations of Chinese, English, and multilingualism, respectively.
Feature Training Data (Language) Link
Visual Modality
VideoMAE-base Kinetics-400 huggingface.co/MCG-NJU/videomae-base
VideoMAE-large Kinetics-400 huggingface.co/MCG-NJU/videomae-large
CLIP-base WIT huggingface.co/openai/clip-vit-base-patch32
CLIP-large WIT huggingface.co/openai/clip-vit-large-patch14
EVA-02-base ImageNet-22k huggingface.co/timm/eva02_base_patch14_224.mim_in22k
DINOv2-large LVD-142M huggingface.co/facebook/dinov2-large
VideoMAE-base (VoxCeleb2) VoxCeleb2 github.com/zeroQiaoba/MERTools
VideoMAE-base (MER2023) MER2023 github.com/zeroQiaoba/MERTools
Acoustic Modality
emotion2vec Mix (EN) github.com/ddlBoJack/emotion2vec
Whisper-base Internet (MULTI, mainly EN) huggingface.co/openai/whisper-base
Whisper-large Internet (MULTI, mainly EN) huggingface.co/openai/whisper-large-v2
wav2vec 2.0-base WenetSpeech (CH) huggingface.co/TencentGameMate/chinese-wav2vec2-base
wav2vec 2.0-large WenetSpeech (CH) huggingface.co/TencentGameMate/chinese-wav2vec2-large
HUBERT-base WenetSpeech (CH) huggingface.co/TencentGameMate/chinese-hubert-base
HUBERT-large WenetSpeech (CH) huggingface.co/TencentGameMate/chinese-hubert-large
Lexical Modality
PERT-base EXT Data (CH) huggingface.co/hfl/chinese-pert-base
PERT-large EXT Data (CH) huggingface.co/hfl/chinese-pert-large
LERT-base EXT Data (CH) huggingface.co/hfl/chinese-lert-base
LERT-large EXT Data (CH) huggingface.co/hfl/chinese-lert-large
XLNet-base EXT Data (CH) huggingface.co/hfl/chinese-xlnet-base
MacBERT-base EXT Data (CH) huggingface.co/hfl/chinese-macbert-base
MacBERT-large EXT Data (CH) huggingface.co/hfl/chinese-macbert-large
RoBERTa-base EXT Data (CH) huggingface.co/hfl/chinese-roberta-wwm-ext
RoBERTa-large EXT Data (CH) huggingface.co/hfl/chinese-roberta-wwm-ext-large
ELECTRA-base EXT Data (CH) huggingface.co/hfl/chinese-electra-180g-base-discriminator
ELECTRA-large EXT Data (CH) huggingface.co/hfl/chinese-electra-180g-large-discriminator
BLOOM-7B ROOTS (MULTI) huggingface.co/bigscience/bloom-7b1
MOSS-7B Mix (MULTI, mainly EN and CH) huggingface.co/fnlp/moss-base-7b
ChatGLM2-6B Mix (MULTI, mainly EN and CH) huggingface.co/THUDM/chatglm2-6b
Baichuan-13B Mix (MULTI, mainly EN and CH) huggingface.co/baichuan-inc/Baichuan-13B-Base